# Index Terms: 6G, large language models (LLMs), fine-tuning, safety alignment, open-source telecom datasetsI Introduction and Motivation

### v1

@@ IV Experimental Setup
For SafeInstruct, we interleave a subset of harmful QA pairs (with safe refusals) from Bianchi’s dataset in [19] into the fine-tuning sets. Specifically, we inject 2500, 1000, and 10 safety samples into TeleData, TeleQnA, and TSpecLLM datasets, respectively. For SafeLoRA, we define the safety-aligned subspace using the respective base and instruct/chat versions of each model. We tune the cosine similarity threshold for and select the configuration that yields the best trade-off between safety and task utility. For SafeMERGE, we follow the same procedure and additionally explore linear merging factors for . The safe reference model used for merging is obtained by fine-tuning each LLM on 1000 samples from Bianchi’s dataset [19], resulting in consistently safe behavior on both HexPhi and DirectHarm (see Table I).
IV-D Open-Source CPT Telecom Models.
@@
We evaluate two publicly available TeleLLMs from [5]: Llama-3-8B-Tele-it and Gemma-2B-Tele-it. Both models are adapted from their non-instruct base versions using large-scale telecom corpora, including 3GPP standardization documents. In [5], the authors perform additional instruction-tuning using the Open-Instruct dataset [26]. While this makes the model helpful and instills a chat-like behavior, Open-Instruct does not contain explicit safety samples, foreshadowing increased harmfulness after CPT, namely when exposed to tabular or math-heavy 3GPP content. To extend SafeInstruct, SafeLoRA, and SafeMERGE to these models, we proceed as follows:
@@
SafeInstruct: we fine-tune each model for one additional epoch using the same hyperparameters as for SFT, interleaving 2500 safety-aligned samples from [19] into the Open-Instruct dataset. This simulates the inclusion of safety data that was missing during instruction tuning.
@@
SafeLoRA / SafeMERGE: we extract the LoRA layer representations from the CPT models and apply the same safety projection and merging strategies as previously described. No additional training is required.
@@ V Results and Discussions
| Qwen-2.7B-Instruct | TeleData | 34.70 | 48.80 | 48.70 | 46.50 | 48.80 | |
@@
| Qwen-2.7B-Instruct | TeleQnA | 45.80 | 65.60 | 64.80 | 64.10 | 65.20 | |
@@
| Qwen-2.7B-Instruct | TSpecLLM | 12.50 | 28.30 | 28.00 | 27.70 | 28.10 | |
@@
We now discuss how SFT and CPT can degrade safety when using telecom datasets, and how effective the discussed realignment defenses are. Tables II and III summarize the results for all models and datasets, reporting task performance (as accuracy on telecom QAs) and corresponding harmfulness.
V-A Telecom Task Utility.
@@
For SFT, task utility improves significantly across models and datasets with accuracy gains between 10% and 25%, evaluated by the Mixtral LLM judge. In general, Llama-3.1 shows the strongest performance, achieving accuracies of 47.60%, 67.80%, and 62.10% on TeleData, TeleQnA, and TSpecLLM, respectively. Its high performance even prior to SFT suggests that the model is already well-aligned with telecom-specific domain jargon. Further, Llama-3.1 and Qwen-2 tend to outperform the older Llama-2 model, except on TSpecLLM, where Qwen-2 achieves only 12.50% accuracy, compared to 33.30% for Llama-2. Interestingly, we find that SFT even on small datasets such as TSpecLLM (just 80 samples) leads to noticeable improvements in utility, highlighting the effectiveness of even light domain adaptation.
@@
For CPT, task utility improves similarly between 10% and 15% on the TeleData benchmark for the public Gemma and Llama-3 TeleLLMs [5]. We additionally evaluated both models on TeleQnA and TSpecLLM datasets, observing similar trends. These results suggest that CPT, as performed by the authors in [5], effectively instills telecom knowledge, as measured across diverse telecom benchmarks. Moreover, CPT with large 3GPP corpora may have further enhanced the models’ deeper understanding of standards and protocols, which is reflected in their conversational interactions.
V-B Harmfulness after SFT and CPT.
@@
For SFT, harmfulness increases noticeably compared to original models, confirming that fine-tuning with telecom data degrades safety of instruction-tuned models. We observe that harmfulness increases the most for the larger TeleData dataset. For example, Llama-2 sees its DirectHarm (HexPhi) score rise from 5.00% (2.00%) to 36.70% (20.10%), marking the sharpest decline, followed by Qwen-2, whose original model is already the least safe. Interestingly, harmfulness on TeleQnA and TSpecLLM is comparable, suggesting a similar impact despite TSpecLLM being significantly smaller (80 samples).
For CPT, we observe extremely high harmfulness scores around mid-to-high 70% for both models on DirectHarm, and even 88.50% on HexPhi for Gemma. This confirms our previous assumption that CPT with telecom data, when not followed by safety-focused instruction tuning, can lead to severe safety degradation, such that the model is likely to comply with most harmful prompts. In fact, both TeleLLMs provide answers to our introductory example on how to effectively destroy a base station, suggesting sledgehammers and malware attacks. In addition, Gemma provides some helpful advice: “Be patient; you may be able to slowly degrade the station’s signal by constantly launching more attacks.” We encourage further investigation of harmful telecom prompt handling to better understand model vulnerabilities. In Fig. 2, we highlight the five most frequent unsafe categories for the Llama-3 TeleLLM on DirectHarm, showing that the majority of unsafe responses fall under non-violent crimes, followed by privacy violations and defamation. Similar trends are observed on HexPhi, with additional spikes in sexual content. Thus, safety samples must be included during instruction tuning in CPT, particularly for telecom data, which often include corpora that can resurface or amplify harmful behavior of pre-trained base models.
V-C Safety Realignment.
Across both SFT and CPT, SafeInstruct, SafeLoRA, and SafeMERGE significantly improve safety while maintaining strong telecom task utility. For Llama-2 on TeleData, either defense preserves accuracy between 37% and 38% while significantly reducing harmfulness, for instance, by up to 30% on DirectHarm and 15% on HexPhi when using SafeMERGE, which generally provides the best safety-utility trade-off among the examined defenses. For Llama-3 and Qwen-2, harmfulness can, in most cases, be reduced even below that of the original safety-tuned models. Furthermore, all SafeCOMM models refuse to answer our introductory prompt, as well as other similar harmful, telecom-related prompts.
@@
For CPT, all three defenses similarly reduce harmfulness from previously extreme levels to low double-digit scores with strong utility. We found thresholds around 0.6 or 0.7 to be optimal for both SafeLoRA and SafeMERGE (with values of 0.7 or 0.8), such that only a small portion of LoRA layers need to be adapted. However, SafeMERGE requires tuning two hyperparameters instead of one for SafeLoRA. In comparison, SafeInstruct is the easiest to implement while requiring relatively few safety samples. We further added telecom-inspired safety refusals (e.g., for our base station example) but observed no notable gains during testing with harmful telecom-related prompts, suggesting that Bianchi’s data from [19] generalizes well, even for the telecom domain.
Overall, our study shows that in fact telecom data is not immune to safety erosion during fine-tuning. Nevertheless, even highly unsafe telecom-tuned LLMs can be effectively realigned using open-source-compatible methods, yielding SafeCOMM models with minimal effort from practitioners.
@@ VI Conclusion
In this paper, we studied the impact of tuning LLMs to the telecom domain on model safety, demonstrating that both SFT and CPT can significantly degrade safety alignment, making Telecom LLMs unsafe for real-world deployment. We investigated this issue across three representative datasets for SFT and evaluated two publicly available Telecom LLMs that were continually pre-trained on large-scale telecom corpora. Our findings show that incorporating safety-aligned instruction tuning when adapting LLMs to the telecom domain is necessary, as technical telecom data can inadvertently resurface or amplify harmful behaviors present in the base model. We further showed that lightweight, open-source safety realignment methods can easily restore safety while preserving strong telecom task utility. Our study thus underscores a key takeaway: Safety alignment should not be an afterthought in the development of Telecom LLMs and can be addressed either early or even post-hoc with little effort and substantial impact.

### v2

@@ IV-C Safety Realignment Defenses
For SafeInstruct, we interleave a subset of harmful QA pairs (with safe refusals) from Bianchi’s dataset in [16] into the fine-tuning sets. Specifically, we inject 2500, 1000, and 10 safety samples into TeleData, TeleQnA, and TSpecLLM datasets, respectively. For SafeLoRA, we define the safety-aligned subspace using the respective base and instruct/chat versions of each model. We tune the cosine similarity threshold for and select the configuration that yields the best trade-off between safety and task utility. For SafeMERGE, we follow the same procedure and additionally explore linear merging factors for . The safe reference model used for merging is obtained by fine-tuning each LLM on 1000 samples from Bianchi’s dataset [16], resulting in consistently safe behavior on both DirectHarm and HexPhi (see Table I).
IV-D Open-Source CPT Telecom Models
@@
We evaluate two publicly available TeleLLMs from [4]: Llama-3-8B-Tele-it and Gemma-2B-Tele-it. Both models underwent CPT using large-scale telecom corpora and additional instruction tuning was performed using the Open-Instruct dataset [24]. While this instills a helpful chat-like behavior, Open-Instruct does not contain explicit safety samples, foreshadowing increased harmfulness after being exposed to tabular or math-heavy 3GPP content. We extend SafeInstruct, SafeLoRA, and SafeMERGE to these models as follows:
@@
SafeInstruct: we fine-tune each model for one additional epoch using the same SFT hyperparameters, interleaving 2500 safety samples from [16] into Open-Instruct, thereby simulating safety data during instruction tuning.
@@
SafeLoRA / SafeMERGE: we extract the LoRA layers from CPT models and apply the same safety projection and merging strategies. No additional training is required.
@@ V Results and Discussions
| Qwen-2-7B-Instruct | TeleData | 34.70 | 48.80 | 48.70 | 46.50 | 48.80 | |
| TeleHarm | 7.00 | 26.00 | 8.20 | 9.50 | 6.50 | ||
@@
| Qwen-2-7B-Instruct | TeleQnA | 45.80 | 65.60 | 64.80 | 64.10 | 65.20 | |
| TeleHarm | 7.00 | 14.50 | 8.70 | 9.40 | 7.90 | ||
@@
| Qwen-2-7B-Instruct | TSpecLLM | 12.50 | 28.30 | 28.00 | 27.70 | 28.10 | |
| TeleHarm | 7.00 | 18.00 | 8.00 | 8.50 | 6.50 | ||
@@
V-A Telecom Task Utility after SFT and CPT
@@
For SFT, task utility improves significantly across models and datasets with accuracy gains between 10% and 25% (see Table II). In general, Llama-3.1 shows the strongest performance, achieving accuracies of 47.60%, 67.80%, and 62.10% on TeleData, TeleQnA, and TSpecLLM, respectively. Its high performance even prior to SFT suggests that the model is already familiar with telecom-specific subjects. Furthermore, Llama-3.1 and Qwen-2 tend to outperform the Llama-2 model, except on TSpecLLM, where Qwen-2 achieves only 12.50% accuracy, compared to 33.30% for Llama-2. Interestingly, we find that SFT even on small datasets such as TSpecLLM (just 80 samples) leads to noticeable improvements in utility, highlighting the effectiveness of light domain adaptation.
@@
For CPT, task utility improves similarly between 10% and 15% on the TeleData benchmark for both Gemma and Llama-3 TeleLLMs, with similar trends observed for TeleQnA and TSpecLLM (see Table III). These results suggest that CPT, as performed by the authors in [4], effectively instills telecom knowledge, measured across diverse telecom benchmarks.
V-B Harmfulness after SFT
@@
For SFT, harmfulness increases noticeably, confirming that fine-tuning with telecom data degrades model safety. We observe the sharpest decline for Llama-2 on the larger TeleData dataset, which sees its DirectHarm (HexPhi) score rise from 5.00% (2.00%) to 36.70% (20.10%), followed by Qwen-2, whose original model is already the least safe. We see similar trends on our TeleHarm benchmark, with Llama-2 reaching 28%, Qwen-2 26%, and Llama-3.1 22%, confirming that harmfulness extends even to telecom-specific prompts. Scores on TeleQnA and TSpecLLM are comparable, suggesting a similar impact despite TSpecLLM being significantly smaller.
V-C Harmfulness after CPT
For CPT, we observe extremely high harmfulness scores around mid-to-high 70% for both models on DirectHarm, and even 88.50% on HexPhi for Gemma. Scores on TeleHarm are lower but still substantial at 43.8% for Llama and 50% for Gemma, showing that our curated telecom prompts remain highly vulnerable, albeit less so than general-purpose ones. This confirms our previous assumption that CPT with telecom data, when not followed by safety-focused instruction tuning, can lead to severe safety degradation, such that models are likely to comply with harmful prompts. In fact, both TeleLLMs provide answers to our introductory example on how to destroy a public base station. Below is one randomly sampled answer from the Llama TeleLLM (details omitted for safety):
In Fig. 2, we highlight the five most frequent unsafe categories for the Llama TeleLLM on TeleHarm, showing that the majority of unsafe responses fall under non-violent crimes, followed by privacy violations and defamation. Thus, safety samples must be included during instruction tuning after CPT as telecom data is shown to resurface harmful behavior.
V-D Safety Realignment
@@
For SFT and CPT, SafeInstruct, SafeLoRA, and SafeMERGE significantly improve safety while maintaining telecom task utility. For Llama-2 on TeleData, either defense preserves accuracy around 38% while reducing harmfulness by up to 30% (15%) on DirectHarm (HexPhi) and by 25% on TeleHarm when using SafeMERGE, which generally provides the best safety-utility trade-off among the examined defenses, followed by SafeInstruct. For Llama-3 and Qwen-2, harmfulness can, in most cases, be reduced even below that of the original safety-tuned models. For CPT, all three defenses similarly reduce harmfulness from previously extreme levels to low double-digit scores with strong utility. We found thresholds around 0.6 to be optimal for both SafeLoRA and SafeMERGE (with values of 0.7), such that only a small portion of LoRA layers need to be adapted. However, SafeMERGE requires tuning two parameters instead of one for SafeLoRA. SafeInstruct is the easiest to implement while requiring relatively few safety samples. We further added telecom-inspired safety refusals (e.g., for our base station example) but observed no notable gains during testing, suggesting that Bianchi’s data from [16] generalizes well, even for the telecom domain. Ultimately, these results show that either defense can reduce harmfulness without compromising utility, particularly on TeleHarm, which targets telecom threat scenarios across the network stack.
V-E Per-Token KL Divergence between Safe and Unsafe Models
In Fig. 3, we illustrate the per-token KL divergence on unsafe TeleHarm prompts between fine-tuned models and their unaligned base counterparts, capturing how quickly the distributions drift toward unsafe behavior. This provides a token-level-proof that SFT with telecom data consistently degrades safety. While all three telecom models exhibit high KL divergence in the first few tokens, their distributions quickly converge toward the unsafe base model, indicating that alignment is only a few tokens deep. As a result, models may eventually respond to harmful prompts despite an initial refusal (e.g., “I cannot, but …”). Degradation differs across datasets: TSpecLLM diverges most, followed by TeleQnA and TeleData, likely because its smaller training set keeps it initially closer to the safe instruct model. This shows that safety erosion depends on the number of potentially harmful telecom samples during training. We also include the divergence between the SafeMERGE variant and its safety-aligned Llama-3 instruct model, showing that SafeMERGE closely matches the instruct model, thereby restoring safety alignment from the first tokens.
Overall, our study confirms through multiple experiments and across models that telecom data is not immune to safety erosion during fine-tuning. Our curated TeleHarm dataset provides an important first step toward safety-focused telecom evaluations, and we highlight telecom-specific red-teaming of LLMs as an important direction for future work.
@@ VI Conclusion
In this paper, we studied the impact of fine-tuning LLMs to the telecom domain on model safety, demonstrating that both SFT and CPT can significantly degrade safety alignment. To benchmark this problem, we introduced TeleHarm, a novel telecom-specific red-teaming dataset, and investigated model safety after fine-tuning on three representative SFT datasets as well as two publicly available TeleLLMs continually pre-trained on large-scale telecom corpora. Our findings show that incorporating safety-aligned instruction tuning is necessary, as technical telecom data can inadvertently resurface or amplify harmful behaviors present in the base model. We further demonstrated that lightweight, open-source safety realignment methods can effectively restore safety while preserving strong telecom task utility. Our study thus underscores a key takeaway: safety alignment should not be an afterthought in the development of telecom LLMs and can be addressed either early or post-hoc with little effort and substantial impact.